Write processes which occur on inert electrodes during electrolysis of FeCl3 and Ca(No3)2 acqueos solution

1
Expert's answer
2022-03-30T13:57:04-0400

Electrolysis of FeCl3

Anode: 2Cl- = Cl2(g) + 2e-

Cathode: Fe3+ + 3e- = Fe(s)

Electrolysis of Ca(NO3)2

Anode: 2H2O - 4e- = O2 + 4H+

Cathode: 2H2O + 2e- = H2 + 2OH-